OpenStudy (anonymous):

how do I solve this: 3x+sqrt(x-1)=x+2

OpenStudy (anonymous):

I need to solve for x, then check the answer I tried the quadratic equation but it got super complicated and was wondering if anyone knew a good way to solve this

OpenStudy (anonymous):

here's my work at the moment... sqrt(x-1)=-2x+2 x-1=4x^2+4 x-5=4x^2 -4x^2+x-5=0 -1+/- sqrt(1-80) /-8 (-1+/-sqrt(-79))/-8 -1/8(-1+i sqrt(79) -1/8(7.9i)=-.9875i -1/8(-1-i sqrt(79) -1/8(-9.9i)=1.2375i 3(-.9875i)+sqrt(-.9875i-1)=-.9875i+2

OpenStudy (cwrw238):

you've gone wrong in the second line (-2x + 2)^2 = 4x^2 - 8x + 4

OpenStudy (anonymous):

Thx... was it right to do the quadratic formula, or was there an easier equation to try?

OpenStudy (cwrw238):

so x-1 = 4x^2 - 8x + 4 4x^2 - 9x + 5 = 0

OpenStudy (cwrw238):

- you can factor this one

OpenStudy (anonymous):

Oh I see can you stay for a bit I might need some help still

OpenStudy (anonymous):

Thank you very much

OpenStudy (anonymous):

(x-1)(4x-5)?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

sorry I'm in algebra 2 and I only learned how to factor and put into the standard form. Can you just tell me the gist on solving for x from here? Sorry

OpenStudy (cwrw238):

ok if ab = 0 then either a and b must be 0 here (x-1)=0 or (4x-5)=0

OpenStudy (cwrw238):

* thats a or b

OpenStudy (anonymous):

Ok so the answer is either 1 or 5/4, just need to check one fits into the orginal problem?

OpenStudy (cwrw238):

right - one of these roots might be extraneous

OpenStudy (anonymous):

Thank you very much.

OpenStudy (cwrw238):

yw